### v3

Abstract
Integrating multiple observational studies for meta-analysis has sparked much interest. The presented R package WMAP (Weighted Meta-Analysis with Pseudo-Population) (Guha et al., 2024) addresses a critical gap in the implementation of integrative weighting approaches for multiple observational studies and causal inferences about various groups of subjects, such as disease subtypes. The package features three weighting approaches, each representing a special case of the unified weighting framework introduced by Guha and Li (2024), which includes an extension of inverse probability weights for data integration settings. It performs meta-analysis on user-inputted datasets as follows: (i) it first estimates the propensity scores for study-group combinations, calculates subject balancing weights, and determines the effective sample size (ESS) for a user-specified weighting method; and (ii) it then estimates various features of multiple counterfactual group outcomes, such as group medians and differences in group means for the mRNA expression of eight genes. Additionally, bootstrap variability estimates are provided. Among the implemented weighting methods, we highlight the FLEXible, Optimized, and Realistic (FLEXOR) method, which is specifically designed to maximize the ESS within the unified framework. The use of the software is illustrated through simulations and a multi-site breast cancer case study based on a simulated dataset modeled after real TCGA data from seven medical centers.
Keywords: pseudo-population; retrospective cohort; unconfounded comparison; weighting.
1 Introduction
When analyzing observational studies, balancing covariates is a crucial step for unconfounded causal comparisons of group potential outcomes (Robins and Rotnitzky, 1995; Rubin, 2007). In diverse research areas, the observed or underlying sampling populations of observational studies, in addition to being unbalanced with respect to the group-specific covariates, are invariably very different from the larger natural population of interest. It is therefore necessary to utilize covariate-balancing techniques like weighting or matching (Lunceford and Davidian, 2004). Weighting methods in observational studies with two subject groups rely on the propensity score, and overwhelmingly, utilize inverse probability weights (IPWs) to achieve covariate balance (Rosenbaum and Rubin, 1983; Li and Li, 2019). However, IPW estimators of group differences are often unstable if one of more subjects have extreme PS values (Li and Li, 2019). Consequently, variations of IPWs motivated by truncated subpopulations have been developed (e.g., Crump et al., 2006; Li and Greene, 2013).
Most weighting methods in the literature provide valid inferences for a covariate-balanced pseudo-population that differs from the larger, natural population of interest for which no random samples are available. For example, IPWs correspond to a so-called combined pseudo-population. Overlap weights minimize the asymptotic variance of the weighted average treatment effect for the overlap pseudo-population (Li et al., 2018). For single observational studies with two or more groups, generalized overlap weights minimize the sum of asymptotic variances of weighted estimators of pairwise group differences (Li and Li, 2019). These approaches have the following shortcomings: (i) they are optimal for restricted outcome types and estimands (typically, contrasts of group means) under theoretical conditions that may not be satisfied in practice; furthermore, the study goals may involve very different estimands than the group mean contrasts of counterfactual outcomes (e.g., percentiles, medians, or pairwise correlations of multivariate group responses) and various unplanned estimands; (ii) they are not appropriate for the meta-analysis of multiple observational studies with each study comprising more than two subject groups.
Motivated by these challenges, Guha and Li (2024) extended the propensity score to the multiple propensity score (MPS). Further, they proposed a general family of pseudo-populations and balancing weights that facilitate the integrative analyses and causal inferences of diverse functionals of group-specific potential outcomes. This unified framework generalizes the aforementioned weighting methods to the meta-analysis of multiple studies with multiple groups. Specifically, as explained in Section 2, IPWs and overlap weights (Li et al., 2018; Li and Li, 2019) are extended to the integrative combined (IC) and integrative generalized overlap (IGO) weights, respectively. Furthermore, by optimizing the effective sample size (ESS) of pseudo-populations within this rich family, Guha and Li (2024) invented the so-called FLEXible, Optimized, and Realistic (FLEXOR) weights and studied the properties of these estimand-agnostic and efficient weighted estimators for quantitative, categorical, or multivariate outcomes. The substantial benefits of FLEXOR relative to extensions of existing weighting methods such as IC and IGO are demonstrated using simulated and TCGA cancer datsets in that paper. The approach formulated by Guha and Li (2024) fills a methodological gap by pioneering a principled, estimand-agnostic integrative causal approach capable of accommodating multiple studies with multiple groups, a structure frequently encountered in contemporary integrative research. We have developed an R package, WMAP (Guha et al., 2024), to implement the three weighting methods. To our knowledge, the present R package is the first implementation of such a method.
The paper is structured as follows. Section 2 reviews the weighting methodology and two-stage inference procedure of the FLEXOR approach. Section 3 demonstrates the use of the WMAP package by analyzing the included demo dataset, which mimics the structure of a multi-site breast cancer study from The Cancer Genome Atlas (TCGA). It presents the meta-analysis of retrospective cohorts as a case study, describes the general workflow, illustrates the inferential procedure, and highlights biologically meaningful results. Section 4 conducts simulation studies using the package. An R script named Guha-Xu-Priyam-Li.R is included with the submission to replicate the results presented in the subsequent sections. Section 5 discusses future directions.
2 Methodology
In a large population of interest, suppose there are $K$ groups of patients about whom nothing is known a priori besides the disease prevalences available from registries. The meta-analysis integrates $J$ retrospective cohorts or studies with each study recording the covariates for each subject. For participant $i$, let $S_{i}$ denote their observational study, $Z_{i}$ denote their group, and $\mathbf{X}_{i}\in\mathcal{X}\subset\mathcal{R}^{p}$ denote their covariate vector, and vector of $L$ outcomes by $\mathbf{Y}_{i}\in\mathcal{R}^{L}$. Writing the $z$th group’s counterfactual outcome vector (i.e., the outcome if the $i$th subject had belonged to group $z$) by $\mathbf{Y}_{i}^{(z)}=(Y_{i1}^{(z)},\ldots,Y_{iL}^{(z)})^{\prime}$, the observed outcome is then $\mathbf{Y}_{i}=\mathbf{Y}_{i}^{(Z_{i})}$. We further assume that (a) $J$ and $K$ are not large, (b) a subject belongs to just one observational study, and (c) subjects belonging to all $K$ groups are observed in every study.
If the subject labels are not meaningful, then $(S_{i},Z_{i},\mathbf{X}_{i},\mathbf{Y}_{i})$ may be regarded as i.i.d. samples from an observed population with density $p_{+}[S,Z,\mathbf{X},\mathbf{Y}]$, where $p_{+}[\cdot]$ represents observed population distributions or densities with respect a dominating measure. Generalizing the assumptions of Rubin (2007) and Imbens (2000), we assume: (I) Stable unit treatment value assumption (SUTVA): Conditional on the covariates, the study and group to which a subject belongs has no effect on their potential outcomes, and every version of grouping would lead to the same potential outcomes; (II) Study-specific unconfoundedness: Given study $S_{i}$ and covariate $\mathbf{X}_{i}$, group $Z_{i}$ is statistically independent of $\mathbf{Y}_{i}^{(1)},\ldots,\mathbf{Y}_{i}^{(K)}$, i.e., $p_{+}[\mathbf{Y}^{(z)}\mid S,Z,\mathbf{X}]=p_{+}[\mathbf{Y}^{(z)}\mid S,\mathbf{X}]$; and (III) Positivity: $p_{+}[S=z,Z=z,\mathbf{X}=\mathbf{x}]>0$ for all $(s,z,\mathbf{x})$. Under these conditions, the presented WMAP package performs a two-stage analysis: Stage 1 computes weights to integrate multiple studies, and Stage 2 uses these weights to infer group counterfactual outcomes.
2.1 Stage 1: Outcome-free sample weights
For the purpose of meta-analysis, we generalize the propensity score (e.g., Rosenbaum and Rubin, 1983) to the multiple propensity score (MPS) of study-group memberships belonging to $\Sigma=$$\{1,\ldots,J\}\times\{1,\ldots,K\}.$ For $\mathbf{x}\in\mathcal{X}$, the MPS is defined as
$\delta_{sz}(\mathbf{x})=p_{+}\bigl{[}S=s,Z=z\mid\mathbf{X}=\mathbf{x}\bigr{]}\quad\text{for $(s,z)\in\Sigma$. }$ (1)
In observational studies, the unknown MPS is estimated by regressing the factor variable $(S_{i},Z_{i})$ on $\mathbf{x}_{i}$ for the $N$ subjects using parametric or nonparametric regression techniques. Using MPS, we can derive subject-specific balancing weights that redistribute, and thereby, transform the density of the (unbalanced) observed population to a covariate-balanced pseudo-population (Guha and Li, 2024) in which a patient’s study, group and covariates are mutually independent by design:
$p\bigl{[}S=s,Z=z,\mathbf{X}=\mathbf{x}\bigr{]}=\gamma_{s}\,\theta_{z}\,f_{\boldsymbol{\gamma},\boldsymbol{\theta}}(\mathbf{x}),\quad\text{for $(s,z,\mathbf{x})\in\Sigma\times\mathcal{X}$,}$ (2)
where $p[\cdot]$ denotes distributions or densities with respect to the pseudo-population, and which relies on probability vector $\boldsymbol{\gamma}=(\gamma_{1},\ldots,\gamma_{J})$ quantifying the study relative masses, relative group prevalences $\boldsymbol{\theta}=(\theta_{1},\ldots,\theta_{K})$, and pseudo-population covariate density $f_{\boldsymbol{\gamma},\boldsymbol{\theta}}(\mathbf{x})$. In some investigations, it is possible to specify group prevalence $\boldsymbol{\theta}$ to match the group prevalence of the natural population. However, reasonable choices of $\boldsymbol{\gamma}$ are often unknown because they are primarily determined by the study designs and unknown factors influencing cohort participation. If any component of $\boldsymbol{\gamma}$ or $\boldsymbol{\theta}$ is unknown, we can optimize the pseudo-population in later steps with respect to these quantities.
Let the marginal observed covariate density be denoted by $f_{+}(\mathbf{x})$, $\mathbf{x}\in\mathcal{X}$. Then, there exists a tilting function (e.g., Li et al., 2018), $\eta_{\boldsymbol{\gamma},\boldsymbol{\theta}}$, such that $f_{\boldsymbol{\gamma},\boldsymbol{\theta}}(\mathbf{x})\propto\eta_{\boldsymbol{\gamma},\boldsymbol{\theta}}(\mathbf{x})f_{+}(\mathbf{x})$, implying that $f_{\boldsymbol{\gamma},\boldsymbol{\theta}}(\mathbf{x})=\eta_{\boldsymbol{\gamma},\boldsymbol{\theta}}(\mathbf{x})f_{+}(\mathbf{x})/\mathbb{E}_{+}[\eta_{\boldsymbol{\gamma},\boldsymbol{\theta}}(\mathbf{X})]$, where $\mathbb{E}_{+}(\cdot)$ denotes expectations under the observed distribution in which $\mathbf{X}\sim f_{+}$. This gives an alternative characterization of pseudo-populations relying on tilting functions instead of covariate densities. In particular, with $\mathcal{S}_{J}$ denoting the unit simplex in $\mathcal{R}^{J}$, different $\boldsymbol{\gamma}\in\mathcal{S}_{J}$, $\boldsymbol{\theta}\in\mathcal{S}_{K}$, and tilting function $\eta_{\boldsymbol{\gamma},\boldsymbol{\theta}}$ result in different pseudo-populations belonging to family (2) as natural meta-analytical extensions of existing weighting methods for single studies. For example, equally weighted studies and groups, along with tilting function $\eta_{\boldsymbol{\gamma},\boldsymbol{\theta}}(\mathbf{x})\propto 1$ and $\eta_{\boldsymbol{\gamma},\boldsymbol{\theta}}(\mathbf{x})=1/\sum_{s}\sum_{z}\delta_{sz}^{-1}(\mathbf{x})$, respectively, extends the combined (Li et al., 2018) and generalized overlap (Li and Li, 2019) pseudo-populations to the integrative combined (IC) and integrative generalized overlap (IGO) pseudo-populations, respectively.
For more generality and efficiency, we define a (multi-study) balancing weight as the ratio of pseudo-population and observed population densities (Guha and Li, 2024). For $(s,z,\mathbf{x})\in\Sigma\times\mathcal{X}$, balancing weight
$\rho_{\boldsymbol{\gamma},\boldsymbol{\theta}}(s,z,\mathbf{x})=\frac{\gamma_{s}\,\theta_{z}\,\eta_{\boldsymbol{\gamma},\boldsymbol{\theta}}(\mathbf{x})}{\delta_{sz}(\mathbf{x})\,\mathbb{E}_{+}[\eta_{\boldsymbol{\gamma},\boldsymbol{\theta}}(\mathbf{X})]}.$ (3)
redistributes the observed distribution’s density to equal the pseudo-population’s density. The unnormalized sample weights, $\tilde{\rho}_{i}=$ $\gamma_{s_{i}}\theta_{z_{i}}\eta_{\boldsymbol{\gamma},\boldsymbol{\theta}}(\mathbf{x}_{i})/\delta_{s_{i}z_{i}}(\mathbf{x}_{i})$, are later used to evaluate unconfounded weighted estimators of various counterfactual outcome features in the pseudo-population. The empirically normalized balancing weight is computed from the $N$ unnormalized sample weights as $\rho_{i}=\tilde{\rho}_{i}/\sum_{u=1}^{N}\tilde{\rho}_{u}$ and is normalized to have sample mean 1.
A widely used, estimand-agnostic measure of a weighting method’s inferential accuracy is the effective sample size (ESS) (e.g., McCaffrey et al., 2013), $\mathcal{Q}(\boldsymbol{\gamma},\boldsymbol{\theta},\eta_{\boldsymbol{\gamma},\boldsymbol{\theta}})=N/\bigl{[}1+\text{Var}_{+}\bigl{\{}\rho_{\boldsymbol{\gamma},\boldsymbol{\theta}}(S,Z,\mathbf{X})\bigr{\}}\bigr{]}=N/\mathbb{E}_{+}\bigl{\{}\rho_{\boldsymbol{\gamma},\boldsymbol{\theta}}^{2}(S,Z,\mathbf{X})\bigr{\}}$, where $\mathbb{E}[\cdot]$ and $\text{Var}[\cdot]$ respectively denote expectations and variances with respect to the pseudo-population. When $N$ is large, the ESS is reliably estimated by the sample ESS, $\tilde{\mathcal{Q}}(\boldsymbol{\gamma},\boldsymbol{\theta},\eta_{\boldsymbol{\gamma},\boldsymbol{\theta}})=N^{2}/\sum_{i=1}^{N}\rho_{i}^{2}$, using the empirically normalized weights.
We define the FLEXOR pseudo-population as a member of family (2) maximizing ESS subject to any investigator-imposed restrictions that $\boldsymbol{\gamma}$ and $\boldsymbol{\theta}$ must belong to $\Upsilon\subset\mathcal{S}_{J}\times\mathcal{S}_{K}$ (Guha and Li, 2024). If the FLEXOR pseudo-population is identified by $(\breve{\boldsymbol{\gamma}},\breve{\boldsymbol{\theta}},\breve{\eta}_{\breve{\boldsymbol{\gamma}},\breve{\boldsymbol{\theta}}})$, then
$\mathcal{Q}(\breve{\boldsymbol{\gamma}},\breve{\boldsymbol{\theta}},\breve{\eta}_{\breve{\boldsymbol{\gamma}},\breve{\boldsymbol{\theta}}})=\sup_{(\boldsymbol{\gamma},\boldsymbol{\theta})\in\Upsilon}\sup_{\eta_{\boldsymbol{\gamma},\boldsymbol{\theta}}}\mathcal{Q}(\boldsymbol{\gamma},\boldsymbol{\theta},\eta_{\boldsymbol{\gamma},\boldsymbol{\theta}}).$
An iterative procedure for estimating FLEXOR pseudo-population
Initializing $(\boldsymbol{\gamma},\boldsymbol{\theta})\in\Upsilon$, we perform the following steps iteratively until the sample ESS converges. The converged pseudo-population with optimized sample ESS estimates the FLEXOR pseudo-population.
  • With the parameters $(\boldsymbol{\gamma},\boldsymbol{\theta})$ held fixed, maximize the sample ESS $\tilde{\mathcal{Q}}(\boldsymbol{\gamma},\boldsymbol{\theta},\eta_{\boldsymbol{\gamma},\boldsymbol{\theta}})$ over all tilting functions to obtain the best fixed-$(\boldsymbol{\gamma},\boldsymbol{\theta})$ pseudo-population represented by $(\boldsymbol{\gamma},\boldsymbol{\theta},\breve{\eta}_{\boldsymbol{\gamma},\boldsymbol{\theta}})$. The analytical form of $\breve{\eta}_{\boldsymbol{\gamma},\boldsymbol{\theta}}$ for the theoretical ESS is given by Theorem 1 of Guha and Li (2024):
    $\breve{\eta}_{\boldsymbol{\gamma},\boldsymbol{\theta}}(\mathbf{x})=\biggl{(}\sum_{s=1}^{J}\sum_{z=1}^{K}\frac{\gamma_{s}^{2}\theta_{z}^{2}}{\delta_{sz}(\mathbf{x})}\biggr{)}^{-1},\quad\text{ $\mathbf{x}\in\mathcal{X}$.}$
    This pseudo-population’s balancing weights (3) are uniformly bounded over $(s,z,\mathbf{x})\in\Sigma\times\mathcal{X}$. Tilting function $\breve{\eta}_{\boldsymbol{\gamma},\boldsymbol{\theta}}(\mathbf{x})$ de-emphasizes covariate regions where $\delta_{sz}(\mathbf{x})$ is nearly 0 for some $(s,z)\in\Sigma$. At the same time, it promotes covariate regions where the group propensities, $\delta_{z}=\sum_{s=1}^{K}\delta_{sz}(\mathbf{x})$, match group proportion $\theta_{z}$ for every $z=1,\ldots,K$. Set function $\eta=\breve{\eta}_{\boldsymbol{\gamma},\boldsymbol{\theta}}$.
  • With tilting function $\eta$ held fixed, maximize the sample ESS $\tilde{\mathcal{Q}}(\boldsymbol{\gamma},\boldsymbol{\theta},\eta)$ over all parameters $(\boldsymbol{\gamma},\boldsymbol{\theta})\in\Upsilon$ to obtain the best fixed-$\eta$ pseudo-population represented by $(\tilde{\boldsymbol{\gamma}},\tilde{\boldsymbol{\theta}},\eta)$. The numerical maximization can be performed using implementations of Gauss-Seidel or Jacobi algorithms. Set $(\boldsymbol{\gamma},\boldsymbol{\theta})=(\tilde{\boldsymbol{\gamma}},\tilde{\boldsymbol{\theta}})$.
The WMAP package includes a function, balancing.weights(), that estimates the MPS and calculates the normalized balancing weights of the $N$ subjects and sample ESS for a user-specified pseudo-population. The required pseudo-population is specified by the user through the argument method, which can be “FLEXOR”, “IC”, or “IGO.” If method="FLEXOR", the iterative procedure, as outlined above, is implemented starting from different initial values to estimate the FLEXOR pseudo-population.
2.2 Stage 2: Unconfounded inferences of group counterfactual outcomes
The theoretical properties (e.g., asymptotic variances) of multivariate weighted estimators of wide-ranging group-level features of the outcomes have been thoroughly investigated (Guha and Li, 2024). We summarize here some special applications of the methodology relevant to the WMAP package implementation. Using univariate outcomes ($L=1$), suppose the potential outcome vectors $Y^{(1)},\ldots,$ $Y^{(K)}$ have common support, $\mathcal{Y}\subset\mathcal{R}$. We assume identical conditional distributions: $p[Y^{(z)}\mid S,Z,\mathbf{X}]=p_{+}[Y^{(z)}\mid S,Z,\mathbf{X}]$ for $z=1,\ldots,K$, guaranteeing that the SUTVA, unconfoundedness, and positivity assumptions for the observed population hold for the pseudo-population. Since the pseudo-population has balanced covariates by design, this implies that $p[Y\mid Z=z]=p[Y^{(z)}]$, paving the way for weighted estimators of potential outcome features for the FLEXOR, IGO, and IC pseudo-populations.
Let $\Phi_{1},\ldots,\Phi_{M}$ be real-valued functions having domain $\mathcal{Y}$. We wish to infer pseudo-population means of transformed potential outcomes, $\mathbb{E}[\Phi_{1}(Y^{(z)})],\ldots,\mathbb{E}[\Phi_{M}(Y^{(z)})]$ for $z=1,\ldots,K$. Appropriate choices of $\Phi_{m}$ correspond to pseudo-population inferences about group-specific marginal means, medians, variances, and CDFs of potential outcome components. Equivalently, writing $\mathbf{\Phi}(Y^{(z)})=\bigl{(}\Phi_{1}(Y^{(z)}),\ldots,\Phi_{M}(Y^{(z)})\bigr{)}^{\prime}\in\mathcal{R}^{M}$, let $\boldsymbol{\lambda}^{(z)}=$ $\mathbb{E}[\mathbf{\Phi}(Y^{(z)})]$. For real-valued functions $\psi$ with domain $\mathcal{R}^{M}$, we wish to estimate $\psi(\boldsymbol{\lambda}^{(z)})$.
For example, define $\Phi_{1}(Y^{(z)})=Y^{(z)}$ and $\Phi_{2}(Y^{(z)})=(Y^{(z)})^{2}$. Then $\psi(t_{1},t_{2})=t_{1}$, we obtain the pseudo-population mean in the $z$th group. Similarly, $\psi(t_{1},t_{2})=\sqrt{t_{2}-t_{1}^{2}}$ gives the pseudo-population standard deviation in the $z$th group. For a second example, let $y_{1},\ldots,y_{M}$ be a fine grid of prespecified points in the support $\mathcal{Y}$ and $\Phi_{m}(Y^{(z)})=\mathcal{I}(Y^{(z)}\leq y_{m})$. For $\psi(t_{1},\ldots,t_{M})=t_{m}$, the pseudo-population CDF of $Y^{(z)}$ evaluated at $y_{m}$ equals $\psi(\boldsymbol{\lambda}^{(z)})$. Similarly, for $\psi(t_{1},\ldots,t_{M})=t_{m^{*}}$ where $m^{*}=\arg\min_{m}|t_{m}-0.5|$, the approximate median of $Y_{1}^{(z)}$ is given by $\psi(\boldsymbol{\lambda}^{(z)})$.
Using the normalized weights $\rho_{1},\ldots,\rho_{N}$, a weighted estimator of mean vector $\boldsymbol{\lambda}^{(z)}$ is
$\mathbf{\bar{\Phi}}_{z}=\frac{\sum_{i=1}^{N}\rho_{i}\,\mathbf{\Phi}(Y_{i})\,\mathcal{I}(Z_{i}=z)}{\sum_{i=1}^{N}\rho_{i}\,\mathcal{I}(Z_{i}=z)},$
and $\psi(\mathbf{\bar{\Phi}}_{z})$ is an estimator of $\psi(\boldsymbol{\lambda}^{(z)})$. Theorem 2 of Guha and Li (2024) shows that these weighted estimators are consistent and asymptotically normal. However, since $N$ may not be sufficiently large to justify asymptotic approximations, the WMAP package applies bootstrap methods to estimate standard errors.
The Stage 2 analysis is implemented in WMAP by function causal.estimate(), which first calls function balancing.weights() to perform the Stage 1 analyses. The function then estimates different features of the $K$ counterfactual group outcomes (e.g., medians and group mean differences) for the IC, IGO, or FLEXOR pseudo-populations. The function also evaluates bootstrap-based variability estimates for these features.
3 Meta-analysis of multiple (multi-site) breast cancer studies
To demonstrate the use of the WMAP package, we analyze the built-in demo dataset, which contains simulated data designed to mirror the structure and characteristics of a multi-site breast cancer study from The Cancer Genome Atlas (TCGA). This illustrative dataset mimics the layout and variable types of the original TCGA study, which involved $J=7$ medical centers and $N=450$ patients, divided into $K=2$ groups based on breast cancer subtypes: infiltrating ductal carcinoma (IDC) and infiltrating lobular carcinoma (ILC). The original dataset, available from the GDC Data Portal upon registration (NCI, 2022), includes $p=30$ unbalanced covariates and $L=8$ outcomes representing mRNA expression levels for selected genes (COL9A3, CXCL12, IGF1, ITGA11, IVL, LEF1, PRB2, and SMR3B) known to be relevant in breast cancer research (Christopoulos et al., 2015). Using the demo dataset, we estimate and compare the counterfactual means, standard deviations, and medians between the IDC and ILC groups. The WMAP package thus provides a convenient and accessible tool for exploring causal meta-analysis methods on synthetic datasets that closely resemble real-world cancer studies, even without direct access to the full TCGA resource.
3.1 Data structure
We begin by installing and loading the WMAP package and the example dataset included in the package.
R> library(WMAP)
R> data(demo)
We then examine the contents of the dataset.
  • X: $p=30$ demographic and clinicopathological covariates.
    R> dim(X)
    [1] 450 30
  • Y: Outcome vectors of mRNA expression measurements for the eight targeted genes arranged in a $450\times 8$ matrix.
    R> round(head(Y),4)
    [,1] [,2] [,3] [,4] [,5] [,6] [,7] [,8]
    [1,] 1.2828 -0.1152 -0.3829 -0.3082 -1.1200 1.2068 -0.9472 0.8768
    [2,] -1.1603 1.5377 1.6034 -0.9822 -0.9507 0.3980 -0.9481 -0.2325
    [3,] -0.3815 1.1320 0.9348 -1.2661 1.1733 -0.0956 -0.1138 2.3797
    [4,] -0.3032 0.5999 1.3941 -0.0299 -1.1010 -0.0838 -0.9565 -0.5058
    [5,] 0.4147 -0.3312 -1.7675 0.7626 -1.1300 -0.6157 -0.4613 -0.9950
    [6,] 0.1867 0.9826 1.2274 0.7895 0.2217 -0.5541 -0.9479 -1.0016
  • S: Site labels of patients, representing the seven medical centers.
  • Z: Group labels of patients, representing the two disease subtypes IDC and ILC.
  • groupnames:
    R> groupNames
    [1] "InfiltratingDuctalCarcinoma"
    [2] "InfiltratingLobularCarcinoma"
  • naturalGroupProp: The relative proportions of IDC and ILC subtypes in the larger U.S. population (Wright, 2022; Tran, 2022).
    R> naturalGroupProp
    [1] 0.8888889 0.1111111
Remark: Users can easily utilize the package functions to conduct meta-analyses on their own datasets. The formatting requirements for user-specified datasets are as follows: (a) Vector S, consisting of $N$ factor levels belonging to the set $\{1,\ldots,J\}$, representing the study memberships of the subjects. Each study must have at least 1 subject; (b) Vector Z, consisting of factor levels belonging to the set $\{1,\ldots,K\}$, representing the $N$ group memberships. Each group should contain at least 1 subject; (c) Covariate matrix, X, of dimension $N\times p$ containing $p$ continuous or binary (0/1) measurements, including factor covariates expanded as dummy binary values; (d) Matrix Y of dimension $N\times L$ comprising $L$ containing outcomes; and (e) Probabilty vector, naturalGroupProp, of length $K$ and strictly positive elements, representing the relative group prevalences $\boldsymbol{\theta}$ of the larger natural population. This last user input is necessary only for FLEXOR weights. It should be skipped for IC and IGO weights, which assume $\boldsymbol{\theta}=(1/K,\ldots,1/K)$; if specified, the input is ignored for these weighting methods.
Input1: S, Z, X, method, seed, naturalGroupProp, num.random, gammaMin, gammaMax
$\triangleright$ See Table I for input argument details
Function balancing.weights(Input1):
      Regress (S,Z) on X and evaluate $\hat{\delta}_{s_{1}z_{1}}(\mathbf{x}_{1}),\ldots,\hat{\delta}_{s_{N}z_{N}}(\mathbf{x}_{N})$triangleright Estimate MPS
      if method="IC" then
             for $i=1,\ldots,N$ do
                   $\tilde{\rho}_{i}\leftarrow$ $1/\hat{\delta}_{s_{i}z_{i}}(\mathbf{x}_{i})$triangleright IC unnormalized weights
             end for
       end if
      if method="IGO" then
             for $i=1,\ldots,N$ do
                   $\eta(\mathbf{x}_{i})\leftarrow 1/\sum_{s=1}^{J}\sum_{z=1}^{K}\hat{\delta}_{sz}^{-1}(\mathbf{x}_{i})$triangleright IGO tilting function
                  $\tilde{\rho}_{i}\leftarrow$ $\eta(\mathbf{x}_{i})/\hat{\delta}_{s_{i}z_{i}}(\mathbf{x}_{i})$triangleright IGO unnormalized weights
             end for
       end if
      if method="FLEXOR" then
             $\boldsymbol{\theta}\leftarrow$ naturalGroupProp $\triangleright$ fixed group prevalence
[…]
                  $\boldsymbol{\gamma}^{(t)}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{S}_{J}\cap\bigl{(}[\text{\tt gammaMin},\text{\tt gammaMax}\bigr{)}]^{J}$triangleright random starting point
                  $\boldsymbol{\gamma}^{(t)}_{{\dagger}},\mathcal{Q}^{(t)}\leftarrow$ FLEXOR.2STEP$\bigl{(}\cdots,\boldsymbol{\theta},\{\hat{\delta}_{s_{i}z_{i}}(\mathbf{x}_{i})\}_{i=1}^{N},\boldsymbol{\gamma}^{(t)}\bigr{)}$
                   $\triangleright$ See Algorithm 2
             end for
            $\breve{t}\leftarrow\text{argmax}_{t}\mathcal{Q}^{(t)}$
            $\breve{\boldsymbol{\gamma}}\leftarrow\boldsymbol{\gamma}^{(\breve{t})}$
            $\mathcal{Q}\leftarrow$ OPTIMIZED.ESS$\bigl{(}\breve{\boldsymbol{\gamma}},\boldsymbol{\theta},\{\hat{\delta}_{s_{i}z_{i}}(\mathbf{x}_{i})\}_{i=1}^{N}\bigr{)}$triangleright See Algorithm 2
            for $i=1,\ldots,n$ do
                   $\eta(\mathbf{x}_{i})\leftarrow$ $\biggl{(}\sum_{s=1}^{J}\sum_{z=1}^{K}\frac{\breve{\gamma}_{s}^{2}\theta_{z}^{2}}{\delta_{sz}(\mathbf{x}_{i})}\biggr{)}^{-1}$
                   $\triangleright$ Best fixed-$(\boldsymbol{\gamma},\boldsymbol{\theta})$ tilting function
             end for
            for $i=1,\ldots,N$ do
                   $\tilde{\rho}_{i}\leftarrow$ $\breve{\gamma}_{s_{i}}\theta_{z_{i}}\eta(\mathbf{x}_{i})/\hat{\delta}_{s_{i}z_{i}}(\mathbf{x}_{i})$triangleright FLEXOR unnormalized weights
             end for
            for $i=1,\ldots,N$ do
                   $\rho_{i}=\tilde{\rho}_{i}/\sum_{u=1}^{N}\tilde{\rho}_{u}$triangleright Empirically normalized weights
             end for
       end if
      return Empirically normalized weights, wt.v $\equiv\{\rho_{i}\}_{i=1}^{N}$ and percentESS $\equiv 100\mathcal{Q}/N$; see Table II for output details.
End Function
Algorithm 1 Stage 1 analysis: Empirically normalized balancing weights
Input: S, Z, X, gammaMin, gammaMax, $\boldsymbol{\theta},\{\hat{\delta}_{s_{i}z_{i}}(\mathbf{x}_{i})\}_{i=1}^{N}$, $\boldsymbol{\gamma}^{(t)}$triangleright passed by Algorithm 1
Function FLEXOR.2STEP(Input):
      $\boldsymbol{\gamma}\leftarrow\boldsymbol{\gamma}^{(t)}$; $\mathcal{Q}^{(\text{new})}=0$; exit=FALSE $\triangleright$ Initialize
      $\boldsymbol{\Gamma}\leftarrow\mathcal{S}_{J}\cap\bigl{(}[\text{\tt gammaMin},\text{\tt gammaMax}\bigr{)}]^{J}$triangleright Admissible values of $\boldsymbol{\gamma}$
      while !exit do
             $\mathcal{Q}^{(\text{old})}\leftarrow\mathcal{Q}^{(\text{new})}$
            $\boldsymbol{\gamma}^{{\dagger}}\leftarrow\text{argmax}_{\boldsymbol{\gamma}\in\boldsymbol{\Gamma}}$ OPTIMIZED.ESS$\bigl{(}\boldsymbol{\gamma},\boldsymbol{\theta},\{\hat{\delta}_{s_{i}z_{i}}(\mathbf{x}_{i})\}_{i=1}^{N}\bigr{)}$triangleright function defined below
            $\mathcal{Q}^{(\text{new})}=$ OPTIMIZED.ESS$\bigl{(}\boldsymbol{\gamma}^{{\dagger}},\boldsymbol{\theta},\{\hat{\delta}_{s_{i}z_{i}}(\mathbf{x}_{i})\}_{i=1}^{N}\bigr{)}$
            if $\mathcal{Q}^{(\text{new})}/\mathcal{Q}^{(\text{old})}-1$ is small then
                   exit $\leftarrow$ TRUE
             end if
       end while
      return $\boldsymbol{\gamma}^{{\dagger}},\mathcal{Q}^{(\text{new})}$
End Function
Function Optimized.ESS($\boldsymbol{\gamma},\boldsymbol{\theta},\{\hat{\delta}_{s_{i}z_{i}}(\mathbf{x}_{i})\}_{i=1}^{N}$):
       $\triangleright$ Maximum ESS for pseudo-population parameters $(\boldsymbol{\gamma},\boldsymbol{\theta})$
      for $i=1,\ldots,n$ do
             $\breve{\eta}_{\boldsymbol{\gamma},\boldsymbol{\theta}}(\mathbf{x}_{i})\leftarrow$ $\biggl{(}\sum_{s=1}^{J}\sum_{z=1}^{K}\frac{\gamma_{s}^{2}\theta_{z}^{2}}{\delta_{sz}(\mathbf{x}_{i})}\biggr{)}^{-1}$
             $\triangleright$ Best fixed-$(\boldsymbol{\gamma},\boldsymbol{\theta})$ tilting function
       end for
      return Sample.ESS$\bigl{(}\boldsymbol{\gamma},\boldsymbol{\theta},\breve{\eta},\{\hat{\delta}_{s_{i}z_{i}}(\mathbf{x}_{i})\}_{i=1}^{N}\bigr{)}$triangleright function defined below
End Function
Function Sample.ESS($\boldsymbol{\gamma},\boldsymbol{\theta},\eta,\{\hat{\delta}_{s_{i}z_{i}}(\mathbf{x}_{i})\}_{i=1}^{N}$):
       for $i=1,\ldots,N$ do
             $\tilde{\rho}_{i}\leftarrow$ $\gamma_{s_{i}}\theta_{z_{i}}\eta(\mathbf{x}_{i})/\hat{\delta}_{s_{i}z_{i}}(\mathbf{x}_{i})$triangleright unnormalized weights
       end for
      for $i=1,\ldots,N$ do
             $\rho_{i}=\tilde{\rho}_{i}/\sum_{u=1}^{N}\tilde{\rho}_{u}$triangleright Empirically normalized weights
       end for
      $\tilde{\mathcal{Q}}(\boldsymbol{\gamma},\boldsymbol{\theta},\breve{\eta}_{\boldsymbol{\gamma},\boldsymbol{\theta}})\leftarrow N^{2}/\sum_{i=1}^{N}\rho_{i}^{2}$
      return $\tilde{\mathcal{Q}}(\boldsymbol{\gamma},\boldsymbol{\theta},\eta_{\boldsymbol{\gamma},\boldsymbol{\theta}})$
End Function
Algorithm 2 Iterative procedure of estimating FLEXOR pseudo-population
Table I: Input arguments for function balancing.weights().
Argument Short description Default
S Vector of factor levels representing the $N$ study -
memberships. Takes values in $\{1,\ldots,J\}$
Z Vector of factor levels representing the $N$
group memberships. Takes values in $\{1,\ldots,K\}$
X Covariate matrix of $N$ rows and $p$ columns -
method Pseudo-population, i.e., weighting method; -
Can be "FLEXOR", "IC", or "IGO"
seed Seed for random number generation NULL
Relevant only when method="FLEXOR"; inputs ignored otherwise
naturalGroupProp Relevant only for FLEXOR pseudo-populations: -
fixed user-specific probability vector $\boldsymbol{\theta}$
num.random Number of random starting points of $\boldsymbol{\gamma}$ 40
in the iterative procedure
gammaMin Lower bound for each $\gamma_{s}$ in the iterative procedure 0.001
gammaMax Upper bound for each $\gamma_{s}$ in the iterative procedure 0.999
Table II: Output list components of function balancing.weights().
Position Names Short description
1 wt.v $N$ empirically normalized sample weights
2 percentESS Percentage sample ESS for pseudo-population
3.2 Workflow of analysis
3.2.1 Stage 1: balancing.weights()
For a prespecified pseudo-population, function balancing.weights() first estimates the MPS and then calculates the subject-specific normalized balancing weights and sample ESS. The input arguments are summarized in Table I. The workflow, outlined in Algorithm 1, is based on the iterative Stage 1 procedure described in Section 2 and detailed in Algorithm 2. The function returns a list of items summarized in Table II. If method is “IC” or “IGO,” many arguments of balancing.weights() are not required and any user-provided values are ignored. For instance, $\gamma_{s}=1/J$ and $\theta_{z}=1/K$ are fixed for these pseudo-populations.
For the FLEXOR pseudo-population, (i) the function assumes that the $K$ group proportions $\boldsymbol{\theta}$ are fixed and specified by the user in vector naturalGroupProp, (ii) optional arguments, gammaMin and gammaMax, represent bounds for each element of the FLEXOR study proportions $\breve{\boldsymbol{\gamma}}$. In other words, $\Upsilon=\mathcal{S}_{J}\cap\bigl{(}[\text{\tt gammaMin},\text{\tt gammaMax}\bigr{)}]^{J}\times\{\boldsymbol{\theta}\}$ in the iterative steps to estimate the FLEXOR pseudo-popuation, and (iii) optional argument num.random indicates the number of random starting points for $(\boldsymbol{\gamma},\boldsymbol{\theta})\in\Upsilon$. The sample ESS maximized over these num.random independent replications identify the estimated FLEXOR pseudo-population, for which the $N$ sample weights and ESS comprise the function’s output.
For example, to calculate the FLEXOR weights, we load the package and data, set a random seed to ensure reproducibility, and set the number of starting points for the iterative procedure before calling balancing.weights():
R> library(WMAP)
R> data(demo)
R> set.seed(1)
R> num.random=25
R> output1 = balancing.weights(S, Z, X, method="FLEXOR",
+ naturalGroupProp=naturalGroupProp, num.random)
FLEXOR... estimate 10
FLEXOR... estimate 20
The output output1 is a result S3 list object of class ‘balancing_weights’, which contains:
  • wt.v: the weights for each individual.
    R> length(output1\$wt.v)
    [1] 450
  • percentESS: the ESS of the FLEXOR weights.
    R> output1\$percentESS
    [1] 34.62166
3.2.2 Stage 2: causal.estimate()
For a prespecified pseudo-population, the function causal.estimate() first calculates the subject-specific normalized balancing weights and sample ESS by a call to the balancing.weights() function. Next, it estimates the means, standard deviations and medians of the counterfactual outcomes of the group $K$, in addition to the mean differences in the group. Finally, the function regenerates the bootstrap samples and estimates the same set of counterfactual features for the bootstrap samples. The input arguments are summarized in Table III. The workflow is illustrated for counterfactual means and SD in Algorithm 3. The function returns a list of items summarized in Table IV.
Inputs2: S, Z, X, Y, B, method, seed, naturalGroupProp, num.random, gammaMin, gammaMax
$\triangleright$ See Table III for input argument details
Function Causal.Estimate(Inputs2):
[…]
      wt.v, percentESS $\leftarrow$ BALANCING.WEIGHTS(Inputs1) $\triangleright$ see Algorithm 1
      $\rho_{1},\ldots,\rho_{n}\equiv$ wt.v $\triangleright$ empirically normalized weights
      for $z=1,\ldots,K$ do
             $\hat{\lambda}_{z}\leftarrow\frac{\sum_{i=1}^{N}\rho_{i}\,Y_{i}\,\mathcal{I}(Z_{i}=z)}{\sum_{i=1}^{N}\rho_{i}\,\mathcal{I}(Z_{i}=z)}$triangleright estimated $z$th group’s counterfactual mean
            $\hat{\sigma}_{z}\leftarrow\biggl{(}\frac{\sum_{i=1}^{N}\rho_{i}\,Y_{i}^{2}\,\mathcal{I}(Z_{i}=z)}{\sum_{i=1}^{N}\rho_{i}\,\mathcal{I}(Z_{i}=z)}-\hat{\lambda}_{z}^{2}\biggr{)}^{1/2}$triangleright estimated $z$th group’s counterfactual SD
       end for
      for $b=1,\ldots,B$ do
             Draw bootstrap sample Sb, Zb, Xb, Yb
[…]
            wt_b.v, percentESS_b $\leftarrow$ BALANCING.WEIGHTS(Inputs1b) $\triangleright$ see Algorithm 1
            $\rho_{1b},\ldots,\rho_{nb}\equiv$ wt_b.v $\triangleright$ empirically normalized weights
            for $z=1,\ldots,K$ do
                   $\hat{\lambda}_{zb}\leftarrow\frac{\sum_{i=1}^{N}\rho_{ib}\,Y_{ib}\,\mathcal{I}(Z_{ib}=z)}{\sum_{i=1}^{N}\rho_{ib}\,\mathcal{I}(Z_{ib}=z)}$triangleright estimated $z$th group’s counterfactual mean
                  $\hat{\sigma}_{zb}\leftarrow\biggl{(}\frac{\sum_{i=1}^{N}\rho_{ib}\,Y_{ib}^{2}\,\mathcal{I}(Z_{ib}=z)}{\sum_{i=1}^{N}\rho_{ib}\,\mathcal{I}(Z_{ib}=z)}-\hat{\lambda}_{zb}^{2}\biggr{)}^{1/2}$triangleright estimated $z$th group’s counterfactual SD
             end for
       end for
      return percentESS, estimates of means $\{\hat{\lambda}_{z}\}_{z=1}^{K}$, SDs $\{\hat{\sigma}_{z}\}_{z=1}^{K}$, bootstrap means $\{\hat{\lambda}_{zb}\}_{z=1}^{K}$, bootstrap SDs $\{\hat{\sigma}_{zb}\}_{z=1}^{K}$. See Table II for all outputs of the actual implementation.
End Function
Algorithm 3 Causal estimation of counterfactual means and SDs
Using the example dataset included with the WMAP package, we provide a step-by-step guide to the causal estimation of different features of the group-specific counterfactual outcomes for the three weighting methods, FLEXOR, integrative combined (IC), and integrative generalized overlap (IGO). As mentioned, the $K=2$ groups of the example dataset simulate the breast cancer subtypes, IDC and ILC, and the $L=8$ outcomes corresponding to the mRNA expression levels of the targeted breast cancer genes, COL9A3, CXCL12, IGF1, ITGA11, IVL, LEF1, PRB2, and SMR3B. The goal is unconfounded, covariate-balanced inference about the group counterfactual means, standard deviations, and medians, as well as counterfactual differences in group means and ratios of group standard deviations.
If necessary, load the necessary packages and data, set a random seed, and set the runtime parameters:
R> library(WMAP)
R> data(demo)
R> set.seed(1)
R> num.random=25
R> B=200
Table III: Input arguments for function causal.estimate().
Argument Short description Default
S Vector of factor levels representing the $N$ study -
memberships. Takes values in $\{1,\ldots,J\}$
Z Vector of factor levels representing the $N$
group memberships. Takes values in $\{1,\ldots,K\}$
X Covariate matrix of $N$ rows and $p$ columns -
Y Matrix of $L$ outcomes, dimension $N\times L$
B Number of bootstrap samples for variance estimation 100
method Pseudo-population, i.e., weighting method; -
Can be "FLEXOR", "IC", or "IGO"
seed Seed for random number generation NULL
Relevant only when method="FLEXOR"; inputs ignored otherwise
naturalGroupProp Relevant only for FLEXOR pseudo-populations: -
fixed user-specific probability vector $\boldsymbol{\theta}$
num.random Number of random starting points of $\boldsymbol{\gamma}$ 40
in the iterative procedure
gammaMin Lower bound for each $\gamma_{s}$ in the iterative procedure 0.001
gammaMax Upper bound for each $\gamma_{s}$ in the iterative procedure 0.999
Then, call causal.estimate() setting method equal to "FLEXOR", "IGO", or "IC". For example, the following command applies the FLEXOR weighting method.
R> output2.f = causal.estimate(S, Z, X, Y, B, method="FLEXOR",
+ naturalGroupProp=naturalGroupProp, num.random)
Table IV: Output list components of function causal.estimate().
Position Names Short description
1 percentESS Percentage sample ESS of pseudo-population
2 moments.ar Array of dimension $3\times K\times L$, containing
$\bullet$ estimated means, SDs, and medians (dimension 1)
$\bullet$ for $K$ groups (dimension 2)
$\bullet$ and $L$ counterfactual outcomes (dimension 3)
3 otherFeatures.v Estimated mean group differences for $L$ outcomes
4 collatedMoments.ar Array of dimension $3\times K\times L\times B$, containing
$\bullet$ moments.ar of $b$th bootstrap sample (dimensions 1–3)
$\bullet$ for $B$ bootstrap samples (dimension 4)
5 collatedOtherFeatures.mt Matrix of dimension $L\times B$ containing
$\bullet$ otherFeatures.v of $b$th bootstrap sample (dimension 1)
$\bullet$ for $B$ bootstrap samples (dimension 2)
6 collatedESS A vector of length $B$, containing
$\bullet$ percentage sample ESS for $B$ bootstrap samples
7 method Pseudo-population method, i.e., weighting method.
The output output2.f is a result S3 list object of class ‘causal_estimates’, which contains:
  • percentESS: the ESS of the FLEXOR weights.
    R> output2.f\$percentESS
    [1] 34.62166
  • moments.ar: the means, standard deviations, and medians of the mRNA expression of the 8 genes in the $K=2$ groups.
    R> output2.f\$moments.ar
    , , Y 1
    group 1 group 2
    mean -0.06417815 -0.08867675
    sd 0.92858998 0.60111155
    median -0.11474907 -0.05937943
    , , Y 2
    group 1 group 2
    mean 0.006908613 0.3910141
    sd 0.980184390 0.8207180
    median 0.092934721 0.4674662
    , , Y 8
    group 1 group 2
    mean -0.6221647 0.08445272
    sd 0.7286661 1.02987578
    median -0.8699967 -0.10213382
  • otherFeatures.v: the mean differences of the 8 genes between the two groups.
    R> output2.f\$otherFeatures.v
    Y 1 Y 2 Y 3 Y 4
    0.0244986 -0.3841055 -0.6304299 0.2305657
    Y 5 Y 6 Y 7 Y 8
    0.4334559 -0.2463727 -0.1624729 -0.7066175
  • collatedMoments.ar: the moments.ar for each bootstrap.
  • collatedOtherFeatures.mt: the mean differences of the 8 genes between the two groups (otherFeatures.v) for each bootstrap sample.
Based on the bootstrap results, we can calculate 95% confidence intervals, for example, for the mean differences of the eight genes:
R> CI.f = round(t(apply(output2.f\$collatedOtherFeatures.mt, 1,
+ function(x) quantile(x,probs = c(0.025,0.975)))),2)
R> CI.f
2.5% 97.5%
Y 1 -0.42 0.27
Y 2 -0.99 -0.24
Y 3 -1.12 -0.38
Y 4 -0.49 0.36
Y 5 0.03 0.66
Y 6 -0.67 0.00
Y 7 -0.21 0.20
Y 8 -1.11 -0.24
To calculate the 95% confidence intervals of the mean, median, and standard deviation for the mRNA expression levels of the eight genes in the two groups, we implement the following:
R> f.moments.ci = apply(output2.f\$collatedMoments.ar, c(1, 2, 3), function(x) {
+ quantile(x, probs = c(0.025,0.975))
R> f.moments.ci
, , group 1, Y 1
mean sd median
2.5% -0.2313879 0.8560743 -0.30739645
97.5% 0.2133786 1.2549474 0.07666521
, , group 2, Y 1
mean sd median
2.5% -0.2179196 0.4631803 -0.2275546
97.5% 0.3311190 0.8063285 0.4274931
, , group 1, Y 8
mean sd median
2.5% -0.7730069 0.4080909 -0.9080523
97.5% -0.3660962 1.1670714 -0.8162318
, , group 2, Y 8
mean sd median
2.5% -0.3272660 0.8267197 -0.7755716
97.5% 0.4901018 1.2365001 0.5687652
To include the 95% CI in the output, we define a function write_res:
R> write_res = function(estimates, CI){
+ lower_bound <- CI[, 1] # 2.5% confidence bound
+ upper_bound <- CI[, 2] # 97.5% confidence bound
+ sapply(1:length(estimates), function(i) {
+ paste0(round(estimates[i],2), "(", round(lower_bound[i], 2), ",",
+ round(upper_bound[i], 2), ")")
and then use write_res to output results including point estimates and CIs separately for each comparison group:
R> f.moments = list()
R> for(i in 1:8){
+ f.moments[[i]] = cbind(group1 = write_res(output2.f\$moments.ar[,1,i],
+ t(f.moments.ci[,,1,i])), # group 1 mean sd median
+ group2 = write_res(output2.f\$moments.ar[,2,i],
+ t(f.moments.ci[,,2,i]))) # group 2 mean sd median
R> f.moments
[[1]]
group1 group2
[1,] "-0.06(-0.23,0.21)" "-0.09(-0.22,0.33)"
[2,] "0.93(0.86,1.25)" "0.6(0.46,0.81)"
[3,] "-0.11(-0.31,0.08)" "-0.06(-0.23,0.43)"
[[2]]
group1 group2
[1,] "0.01(-0.15,0.27)" "0.39(0.29,0.97)"
[2,] "0.98(0.81,1.11)" "0.82(0.6,1.12)"
[3,] "0.09(-0.12,0.32)" "0.47(0.39,1.13)"
[[8]]
group1 group2
[1,] "-0.62(-0.77,-0.37)" "0.08(-0.33,0.49)"
[2,] "0.73(0.41,1.17)" "1.03(0.83,1.24)"
[3,] "-0.87(-0.91,-0.82)" "-0.1(-0.78,0.57)"
For the other implemented weighting methods, i.e., IGO and IC, we would change the method argument in causal.estimate() to "IGO" and "IC", and follow the same procedures as above to obtain estimates and confidence intervals. More specifically, we apply the following commands:
R> output2.igo = causal.estimate(S, Z, X, Y, B, method="IGO")
R> output2.ic = causal.estimate(S, Z, X, Y, B, method="IC")
3.3 Discussion of results
We used causal.estimate() to compute point estimates and 95% confidence intervals, and we summarize the results in Table LABEL:tab1. The analysis was conducted on a simulated dataset modeled after a multi-site TCGA breast cancer study. While the dataset is not based on actual patient data, it reflects key structural and biological features commonly found in real-world genomic studies. This setup allows us to assess the performance of the methods under controlled conditions.
In this simulated setting, all three methods (FLEXOR, IC, and IGO) consistently indicated no significant difference in the counterfactual mean expression of the COL9A3 gene between the invasive ductal carcinoma (IDC) and invasive lobular carcinoma (ILC) groups. However, FLEXOR identified significantly greater variability in expression for IDC, which is consistent with the known biological heterogeneity of this subtype in real datasets (Wang et al., 2024). This kind of variability could be important for understanding differential treatment responses or disease progression.
The genes CXCL12 and IGF1 showed lower counterfactual mean expression in IDC compared to ILC. These findings align with the biological functions of these genes; CXCL12 is involved in cell migration and metastasis, while IGF1 plays a role in cell growth and survival. Although these patterns were observed in simulated data, they demonstrate how meta-causal analysis can help detect group-level trends in gene expression that may be relevant in real studies (Vanden Bempt et al., 2005).
For both CXCL12 and IGF1, the variability in expression did not differ meaningfully between subtypes. This suggests that while the mean expression levels diverged, the underlying mechanisms driving variability may be similar. In actual biological systems, this could reflect regulation through shared pathways or constraints on gene expression stability (Samani et al., 2007).
Although the results are based on simulated data, they highlight how the methods in the WMAP package can support causal interpretation of group differences in gene expression, including both average levels and variability. These types of analyses may have practical applications in identifying subtype-specific markers or informing personalized treatment strategies (McCart Reed et al., 2021).
In nearly all scenarios, FLEXOR produced narrower confidence intervals than both IC and IGO. This higher precision illustrates FLEXOR’s strength in practice, particularly when theoretical conditions do not fully hold. By emphasizing stability and practical feasibility in weight estimation, FLEXOR provides a reliable approach to causal meta-analysis in genomic settings.
Table V: For three targeted genes, COL9A3, CXCL12 and IGF1, the estimates and 95% bootstrap confidence levels (shown in parenthesis) of different population-level estimands of the potential outcomes of group 1 (IDC cancer subtype, denoted by superscript 1) and group 2 (ILC cancer subtype, denoted by superscript 2) with FLEXOR, IC, and IGO weights. An IC or IGO confidence interval is bolded if it is wider than the FLEXOR confidence interval. $\lambda$: group mean; $\sigma$: group standard deviation; $M$: group median.
COL9A3
Estimand FLEXOR IC IGO
$\lambda^{(1)}$ $-0.06(-0.23,0.21)$ $0.06\boldsymbol{(-0.31,0.32)}$ $-0.07\boldsymbol{(-0.37,0.36)}$
$\lambda^{(2)}$ $-0.09(-0.22,0.33)$ $-0.03\boldsymbol{(-0.23,0.45)}$ $-0.05\boldsymbol{(-0.25,0.52)}$
$\sigma^{(1)}$ $0.93(0.86,1.25)$ $1.13\boldsymbol{(0.8,1.27)}$ $1.04\boldsymbol{(0.78,1.34)}$
$\sigma^{(2)}$ $0.6(0.46,0.81)$ $0.69\boldsymbol{(0.45,0.86)}$ $0.68\boldsymbol{(0.43,0.86)}$
$M^{(1)}$ $-0.11(-0.31,0.08)$ $-0.12\boldsymbol{(-0.47,0.27)}$ $-0.18\boldsymbol{(-0.48,0.24)}$
$M^{(2)}$ $-0.06(-0.23,0.43)$ $0.06\boldsymbol{(-0.45,0.53)}$ $-0.06\boldsymbol{(-0.42,0.67)}$
$\lambda^{(1)}-\lambda^{(2)}$ $0.02(-0.42,0.17)$ $0.09\boldsymbol{(-0.61,0.42)}$ $-0.01\boldsymbol{(-0.59,0.35)}$
${\sigma^{(1)}}/{\sigma^{(2)}}$ $1.54(1.17,2.31)$ $1.64\boldsymbol{(1.17,2.33)}$ $1.52\boldsymbol{(1.07,2.58)}$
CXCL12
Estimand FLEXOR IC IGO
$\lambda^{(1)}$ $0.01(-0.15,0.27)$ $-0.11\boldsymbol{(-0.25,0.34)}$ $-0.11\boldsymbol{(-0.26,0.35)}$
$\lambda^{(2)}$ $0.39(0.29,0.97)$ $0.52\boldsymbol{(0.2,1.11)}$ $0.55\boldsymbol{(0.3,1.11)}$
$\sigma^{(1)}$ $0.98(0.81,1.11)$ $1.07\boldsymbol{(0.78,1.13)}$ $1\boldsymbol{(0.76,1.2)}$
$\sigma^{(2)}$ $0.82(0.6,1.12)$ $0.81(0.57,1.09)$ $0.83\boldsymbol{(0.51,1.13)}$
$M^{(1)}$ $0.09(-0.12,0.32)$ $0.04\boldsymbol{(-0.31,0.56)}$ $-0.07\boldsymbol{(-0.25,0.44)}$
$M^{(2)}$ $0.47(0.39,1.13)$ $0.68\boldsymbol{(0.07,1.36)}$ $0.72\boldsymbol{(0.16,1.38)}$
$\lambda^{(1)}-\lambda^{(2)}$ $-0.38(-0.99,-0.24)$ $-0.63\boldsymbol{(-1.15,-0.13)}$ $-0.66\boldsymbol{(-1.18,-0.24)}$
${\sigma^{(1)}}/{\sigma^{(2)}}$ $1.19(0.84,1.7)$ $1.32\boldsymbol{(0.79,1.74)}$ $1.21\boldsymbol{(0.73,1.94)}$
IGF1
Estimand FLEXOR IC IGO
$\lambda^{(1)}$ $0.16(-0.1,0.29)$ $0.12\boldsymbol{(-0.2,0.33)}$ $0.15\boldsymbol{(-0.25,0.38)}$
$\lambda^{(2)}$ $0.79(0.55,1.14)$ $0.84\boldsymbol{(0.5,1.23)}$ $0.88\boldsymbol{(0.52,1.24)}$
$\sigma^{(1)}$ $0.93(0.8,1.1)$ $1.02\boldsymbol{(0.77,1.15)}$ $0.98\boldsymbol{(0.75,1.19)}$
$\sigma^{(2)}$ $0.74(0.5,1.07)$ $0.75(0.48,1.05)$ $0.78\boldsymbol{(0.44,1.06)}$
$M^{(1)}$ $0.26(-0.09,0.35)$ $0.32\boldsymbol{(-0.22,0.55)}$ $0.32\boldsymbol{(-0.23,0.5)}$
$M^{(2)}$ $0.91(0.78,1.23)$ $0.96\boldsymbol{(0.57,1.36)}$ $0.96\boldsymbol{(0.65,1.36)}$
$\lambda^{(1)}-\lambda^{(2)}$ $-0.63(-1.12,-0.38)$ $-0.72\boldsymbol{(-1.27,-0.28)}$ $-0.74\boldsymbol{(-1.3,-0.33)}$
${\sigma^{(1)}}/{\sigma^{(2)}}$ $1.26(0.86,1.96)$ $1.36\boldsymbol{(0.92,2.09)}$ $1.25\boldsymbol{(0.87,2.22)}$
4 Simulations
We used the WMAP package to generate simulated datasets for evaluating various weighting strategies used in inferring population-level characteristics across two subject groups. Specifically, we analyzed these datasets using the causal.estimate() function. We briefly describe the generation strategy below and present the results. Following Guha and Li (2024) and in alignment with the motivating TCGA breast cancer studies, we simulated $R=250$ independent datasets, each containing $J=7$ observational studies, $K=2$ groups, and $L=1$ (i.e., univariate) outcome for $\tilde{N}=500$ subjects. The covariate vectors were sampled with replacement from the demo dataset included in the WMAP package, which mimics the structure of the $N=450$ TCGA breast cancer dataset used in the motivating example. To demonstrate the practical application of the WMAP package, we present results from a representative simulation scenario designed to reflect key features of real-world retrospective studies with confounding. This scenario effectively demonstrates the comparative performance of different weighting strategies implemented in the package. For the reported results, we used the procedure in Section 2 to meta-analyze the seven studies within each artificial dataset.
As an initial step for all $250$ artificial datasets, we conducted k-means clustering on the covariates, $\mathbf{X}_{1},\ldots,\mathbf{X}_{N}$, from the WMAP demo dataset to identify lower-dimensional structure, grouping them into $Q=12$ clusters with centers $\mathbf{q}_{1},\ldots,\mathbf{q}_{Q}\in\mathcal{R}^{p}$ and allocated covariate counts $m_{1},\ldots,m_{Q}$. For each artificial dataset $r=1,\ldots,250$, containing $\tilde{N}$ patients, we then generated the data as follows:
  1. 1.
    Natural population  For the $r$th artificial dataset, and using the Dirichlet distribution on simplex $\mathcal{S}_{Q}$, generate the relative masses of the $Q$ clusters, $\boldsymbol{\pi}^{(r)}=(\pi_{1}^{(r)},\ldots,\pi_{Q}^{(r)})\sim\mathcal{D}_{Q}(\mathbf{1}_{Q})$, with $\mathbf{1}_{Q}$ denoting the vector of $Q$ ones. Fix the patient population size of the large natural population as $N_{0}=10^{6}$, and sample their cluster memberships from a mixture distribution on the first $Q$ natural numbers: $c_{ir}^{(0)}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\sum_{u=1}^{Q}\pi_{u}^{(r)}\zeta_{u}$, where $\zeta_{u}$ denotes a point mass at $u$. Select covariate $\mathbf{x}_{ir}^{(0)}$ uniformly from the $m_{c_{ir}^{(0)}}$ TCGA covariates assigned to the $c_{ir}^{(0)}$th k-means cluster (as above).
    Generate the group proportions in the natural population by drawing $\boldsymbol{\theta}^{(r)}\sim\mathcal{D}_{K}(\mathbf{1}_{K})$, for $K=2$ groups. Define the group-covariate relationships: $\delta^{(r)}{z}(\mathbf{x})\propto 1$ if $z=1$ and $\delta^{(r)}{z}(\mathbf{x})\propto\exp\bigl{(}\omega_{0}^{(r)}+\omega_{1}^{(r)}\sum_{t=1}^{p}x_{t}/\frac{1}{N_{0}}\sum_{i=1}^{N_{0}}\sum_{t=1}^{p}x_{irt}^{(0)}\bigr{)}$ if $z=2$. Here, we set $\omega_{1}^{(r)}=1$ and choose $\omega_{0}^{(r)}$ such that the population average of $\delta^{(r)}_{z}(\mathbf{x}_{ir}^{(0)})$ equals $\theta_{z}^{(r)}$.
  2. 2.
    Covariates  For each subject $i=1,\ldots,\tilde{N}$, select their covariate vector $\tilde{\mathbf{x}}_{i}^{(r)}=(\tilde{x}_{i1}^{(r)},\ldots,\tilde{x}_{ip}^{(r)})^{\prime}$ by sampling with replacement from the $N=450$ demo dataset covariate vectors.
  3. 3.
    Study and group memberships  The study assignment $s_{i}^{(r)}$ and group label $z_{i}^{(r)}$ for each individual were generated as follows:
    1. (a)
      MPS  Define the group-specific study propensities as follows:
      $\log\bigl{(}\delta_{S=s\mid Z=z}(\mathbf{x})/\delta_{S=1\mid Z=z}(\mathbf{x})\bigr{)}=sz\omega_{1}^{(r)}\sum_{t=1}^{p}\tilde{x}_{it}^{(r)}/\frac{1}{\tilde{N}}\sum_{i^{\prime}=1}^{\tilde{N}}\sum_{t=1}^{p}\tilde{x}_{i^{\prime}t}^{(r)}$
      for $s=2,\ldots,J$ and $z=1,2$. Assuming the group propensity scores are the same as in the natural population, the marginal propensity score (MPS) is given by $\delta_{sz}(\mathbf{x})=\delta_{s\mid z}(\mathbf{x})\delta_{z}(\mathbf{x})$. For each patient $i=1,\ldots,\tilde{N}$, evaluate their probability vector $\boldsymbol{\delta}^{(r)}(\mathbf{x}_{i})=\bigl{(}\delta_{11}^{(r)}(\mathbf{x}_{i}),\ldots,\delta_{JK}^{(r)}(\mathbf{x}_{i})\bigr{)}$.
    2. (b)
      Study-group memberships  For each patient $i=1,\ldots,\tilde{N}$, generate $(s_{i}^{(r)},z_{i}^{(r)})$ from a categorical distribution with parameter $\boldsymbol{\delta}^{(r)}(\mathbf{x}_{i})$.
  4. 4.
    Subject-specific observed outcomes  Generate
    $Y_{i}^{(r)}\,\Big{|}\,Z_{i}=z_{i}^{(r)}\stackrel{{\scriptstyle\text{indep}}}{{\sim}}\left(z_{i}^{(r)}\sum_{t=1}^{p}\tilde{x}_{it}^{(r)}+50,\,\tau_{r}^{2}\right),$
    where $\tau_{r}^{2}$ is selected to achieve an approximate $R$-squared of $0.9$.
Next, we set aside knowledge of the simulation parameters and analyzed each artificial dataset using the procedure described in Section 2 for the IC, IGO, and FLEXOR pseudo-populations, as implemented in the WMAP package. Define percent ESS as the effective sample size (ESS) scaled for 100 participants. For the 250 simulated datasets, the first panel of Figure 1 displays boxplots of the percent ESS for the FLEXOR, IGO, and IC pseudo-populations. The IC and IGO pseudo-populations showed comparable ESS. The FLEXOR pseudo-population, however, consistently achieved substantially higher ESS across all datasets.
We employed the Stage 2 strategy described in Section 2, which is implemented in the WMAP package, to estimate the average treatment effect (ATE), defined as the difference in group-level counterfactual means, $\lambda^{(1)}-\lambda^{(2)}$. This was achieved by making weighted inferences under each method’s pseudo-population, where covariates were balanced to support causal interpretation. Because each estimator targets the ATE within its respective pseudo-population, we evaluated accuracy by comparing the estimated values to their corresponding true ATEs obtained via Monte Carlo simulation. The second and third panels of Figure 1 show the absolute biases and standard deviations of the FLEXOR, IGO, and IC estimators across 250 synthetic datasets. For each dataset and method, these performance metrics were computed using 200 independent bootstrap samples.
Overall, while the IGO and IC weighting strategies showed comparable performance in estimating the average treatment effect (ATE) across the simulated datasets, FLEXOR consistently outperformed both, as shown in Figure 1. Additionally, it achieved lower absolute bias and standard deviation than the competing methods in all 250 artificial datasets. Notably, although IGO weights are theoretically optimal for estimating the ATE under certain idealized conditions, such as homoscedastic outcomes and correct model specification (see Li and Li, 2019, for single studies), these assumptions did not hold in our simulation setting. The strong empirical performance of FLEXOR in this context underscores its robustness and practical value, particularly when theoretical optimality conditions are violated. By placing greater emphasis on the stability and practical feasibility of the estimated balancing weights, FLEXOR offers a robust and reliable approach to ATE estimation in meta-analytical contexts, especially in scenarios where the assumptions underlying asymptotic efficiency may be less tenable.
Figure 1: Boxplots summarizing the results across 250 simulated datasets ($\tilde{N}=500$ subjects each): (i) percentage effective sample size (ESS), (ii) absolute bias, and (iii) standard deviation of the WATE (Weighted Average Treatment Effect) under three pseudo-populations.
5 Conclusions and Future Developments
Integrating multiple observational studies to make unconfounded causal or descriptive comparisons of group potential outcomes in large natural populations presents significant challenges, because of the complexities involved in data heterogeneity, selection bias, and the need for accurate balancing across datasets. Recently, Guha and Li (2024) introduced a unified weighting framework designed to address these challenges by extending inverse probability weighting techniques for integrative analyses. To translate this theoretical framework into practice, we have developed the R package WMAP. This software tool is specifically designed for the integrative analysis of user-specified datasets and implements three advanced weighting approaches, i.e., IC (Integrative Calibration), IGO (Integrative Generalized Optimization), and FLEXOR (Flexible Optimization of Weights). These methods enhance the capacity to estimate multiple propensity scores, compute balancing weights for subjects, evaluate effective sample sizes, and derive various estimands of counterfactual group outcomes. The package also includes functionality for calculating bootstrap variability estimates, which are essential for quantifying the uncertainty of the results.
In our illustrative application, we used WMAP to analyze simulated gene expression data that mimic differences between two major breast cancer subtypes: invasive ductal carcinoma (IDC) and invasive lobular carcinoma (ILC). Although based on synthetic data, the analysis yielded patterns consistent with known biological characteristics of these subtypes, demonstrating WMAP’s potential to inform scientific insights in causal genomics. The package supports both observational studies and multi-arm randomized controlled trials (RCTs), especially when within-study treatment assignment mechanisms are known. Looking ahead, future updates will extend WMAP’s functionality to accommodate hybrid study designs that integrate data from RCTs and retrospective cohorts. Current limitations, such as challenges in handling high-dimensional biomarker data, are being actively addressed. These enhancements aim to make WMAP an even more versatile tool for integrative causal inference and descriptive analysis across diverse biomedical research settings.
Acknowledgments
This work was supported by the National Institutes of Health under award DMS-1854003 to SG, award CA249096 to YL, and awards CA269398 and CA209414 to SG and YL.